Moller Polarimeter/A1n/d2n

From PolWiki
Jump to: navigation, search

Moller measurement summaries for A1n/d2n

This is the "Cliff notes section on Moller polarimetry runs during the Polarized Helium-3 Running.

Note on sign of polarization [1]

Hall A spin dance Jan. 07 th, 2020 [2]

December 12, 2019

Five pass commissioning. Used preliminary <azz>. Randoms were higher than expected. Need more shielding?. result: [3]

December 19, 2019

Determined the optimum quad settings, set the detector HV and performed the collimator scan pretty quickly. Before the collimator scan the measured polarization was 80%, after the scan it was 84.5% Unpolarized background? Summary:[4], Result: [5] Epics:[6]

January 18, 2020

Before this measurement, timed in randoms so that right arm comes exactly 16 ns after left. [7]. During measurement, replaced 8 ns with 6 ns delay on left arm. Reduced left/right HV by 25 V (1245/1225). Ran with collimator 5 in - this reduced backgrounds a little. Q2/Q3 current changed to 1284.604 A. Bleedthrough check, ~ 10 nA. [8] Settings: [9], BPMs: [10] summary: [11], result: [12]

January 27, 2020

Moller tune check and collimator settings[13].The beam current was 1 uA and IHWP IN. During the first three runs (1385-1387), Hall A was changing their slit and attenuator. During the last 3 runs(1388-1390),had the same slit settings. BPMs:[14], summary [15], result: [16]

February 3, 2020

Spin Dance, summary: [17], result: [18]

February 12, 2020

Some troubles delivering beam. Tried to take the other IHWP state but ended up backing out. summary: [19], result: [20]

FLIP RIGHT (88.0) to FLIP LEFT (-90.6) February 17, 2020 ~10am

February 19, 2020

Energy change in injector starting Feb 18th: [21] Measurements corrected for Hall A bleedthru (57 Hz) summary: [22], result: [23]

February 27, 2020

Changed HVs: [24] Noisy BPMS/readbacks: [25] , Large coin rates with no target, no beam, no collimators. Bleedthru/Scraping notes: [26] summary: [27], result: [28]

March 6th, 2020

Collimator 6 and 7 were scanned. Found 6 was impinging on coin rate. In and Out measured. Hall A Bleedthru (33Hz/4700Hz) summary: [29], results:[30]

March 12th, 2020 Checked collimator 6 (2.95 vs 2.65). summary: [31], result: [32]

August 3rd, 2020 Gain matched L/R. High rates with no target and no collimators to start but operator tuned beam and the rates came down. summary: [33], result: [34]

August 12th, 2020 Day shift operator had a lot of trouble tuning. Then no beam for several hours (storms/RF), came back at 7pm. summary: [35], result: [36]

August 25th, 2020

September 3rd, 2020 Measurement completed in 1/2 shift. :) Bleedthru@ 0.78% Summary:[37], result: [38]

September 11th, 2020

September 16th, 2020

Table of good Moller runs

This will be a log of all "good" Moller runs; ie, our "production" runs. It will include some detailed notes when necessary.

Date	  Run	 Pol (%)   Error (%)	Hwien	Flip       Energy(MeV)     Comment
01/18/20
	  1366	  	   	        -29.64	right      10.3857	   Detector Tune Check
 	  1368	  86.31	   0.98	        -29.64	right      10.3857	   Production IHWP OUT
 	  1369	  85.23	   0.98	        -29.64	right      10.3857	   Production IHWP OUT
	  1370	  86.53	   1.49	        -29.64	right      10.3857	   Production IHWP OUT
	  1371	 -86.05	   0.90	        -29.64	right      10.3857	   Production IHWP IN   
	  1372	 -85.69	   0.99	        -29.64	right      10.3857	   Production IHWP IN
	  1374	 -86.44	   1.06	        -29.64	right      10.3857	   Production IHWP IN
	  1375	 -84.96	   1.79	        -29.64	right      10.3857	   Production IHWP IN
	  1376	 -83.91	   1.55	        -29.64	right      10.3857	   Production IHWP IN
	  1377	 -85.18	   1.60	        -29.64	right      10.3857	   Production IHWP IN
	  1378	 -85.74	   0.96	        -29.64	right      10.3857	   Production IHWP IN


Table of Bleed Thru Runs

Date     Run     Comment
1/18     1373     40 Hz Coin (5 kHz at 0.8 uA)

Table of Tune Check Runs

Date     Run     Comment
1/18     1364     1294.6 
1/18     1365     1289.6
1/18     1366     1284.6